User Tools

Site Tools


planning

Introduction

This site is a temporary location to facilitate coordination of smaller experiments at the Soudan Underground Lab after the Shaft Fire. By prioritizing our needs here, we can then present an organized set of jobs to help the crew. This will minimize the flow of emails directly to personnel currently engaged in the recovery. In addition, I will continue to update information (see the “Update” section at the bottom), so that you will know when entry will be possible.

Experimental Needs

Please feel free to login (first time will create account) and edit your section. I have put a few placeholders to show how this might work. Immediate needs refers to jobs you would like done by the mine crew as soon as there is any access. I have already communicated the need for SOLO, Gopher, and CoGeNT to keep radon plateout to a minimum.

update Monday, April 18 – Full Power was restored to the lab over the weekend. Cleaning crews and fiber optic cable splicing will occur through this week. Computers will become available for remote access by the end of the week. Currently expecting access may be as soon as mid-May, but we need to properly schedule visits, so we will being communicating directly with the experiments about their visits near the end of April.

CoGeNT

Immediate needs

Nitrogen purge for radon reduction (LN boiloff or N2 compressed gas)

When switching off power to the experiment, please preserve an AC ground (i.e., do not allow CoGeNT electronics to “float”). Make sure both NIM bins are OFF, same for Canberra and Ortec HV power supplies on top bin, besides all other measures to keep power off the experiment.

Plan for startup

As soon as is possible, please flush room-temperature gas (preferably compressed nitrogen, but any other *moisture-free* gas will do) through the input to the detector Dewar, presently attached to a 200 liter LN2 tank. Do so (at low flow, over one or two days) until the temperature of the gas coming through the latex hose exhaust is measured to be room temperature. Please do not alter (at all) the position of the thermocouple inserted into that exhaust when measuring this temperature. Because the detector cold finger does not reach the bottom of the detector Dewar, and the Dewar is new and carefully sealed, there is not much of a heat load to the bottom ~1/3 of the LN2 presently in there. If this procedure is not performed, that bottom layer could stay there for a number of weeks, keeping the detector from truly thermalizing at room temperature, a condition that must be ensured for a number of days prior to refilling/restarting. Please let us know when this is attempted and how long it took to reach room-T at the exhaust.

Please contact us about ten days before the date when you envision being able to receive one of us in the lab to restart the experiment, also foreseeing continued LN2 availability (must have that before we can think of restarting). We'll have a few more instructions for preparation then. THANKS.

SOLO

Immediate needs

Nitrogen purge for radon reduction (LN boiloff or N2 compressed gas)

Plan for startup

  • Require pictures of the area to show debris
  • Determine if vacuum space requires repumping. Where is the vacuum port? Inside or outside of shield?
    • If outside, definitely pump
  • Proceed after lab has been cleaned
  • Wipe down Mylar
  • Restart LN, bias, take initial background spectrum
    • If this is OK, proceed with regular operation
  • If background shows contamination, open chamber and wipe down; open Mylar shield and clean outer Pb layers, inner Mylar surface
  • Retake background spectrum
  • If background still shows contamination, full disassembly and cleaning is likely necessary

Gopher

Immediate needs

Nitrogen purge for radon reduction (LN boiloff or N2 compressed gas)

Plan for startup

Neutron Multiplicity Meter

Immediate needs

No immediate needs… the detector should have weathered the loss in power relatively unharmed. We won't know for sure, however, until the power returns. Continued updates regarding the status of the cleanup and the return of power are our only short-term needs… various individuals have and are providing adequate information.

Plan for startup

Once we have power, I will need some on-site assistance in turning our DAQ computers & electronics back on. We will need to assess the state of the three UPS backups… provided the backups can still provide power to their respective components, the experiment should be operable again almost immediately. Again, I'll know more once we have power. -Ray Bunker (805.280.1076 … raybunker@gmail.com)

Veto Shield

Immediate needs

Make sure that the HV is all off while power is intermittant.

We would like to have the 5 veto computers up and running as soon as possible, so we can continue our DAQ work. The question is when will the network be up, so we can use them remotely? This is a high priority item for us.

Once there is network functionality, a minimum plan would be turning on the veto server and the SE computer at a designated time for a 4-hour window for us to

  • deactivate the LabView License
  • transfer data and software files to TwinCities
  • debug the ADC library software for the HV remote control & readback in order to duplicate it here

Other jobs which could be done in the first week with minimal power are

  • order Ar, CO2 gas for the shield
  • order more hydrogen for the mixer - it is getting very low

Plan for startup

When there is power for the pumps, then the gas rack and gas mixer should be turned on and tested. We can do it remotely provided we can connect to the computers.

Reeves Electroforming

Immediate needs

Plan for startup

Gravity Meter

Immediate needs

=== Plan for startup === The start up has been accomplished by Jerry Meier. A replacement UPS power supply is being shipped to the Surface Building in Soudan for installation in the gravity meter setup.

Update from Marvin Marshak (Saturday, April 16)

• Full electrical power was restored to the entire Soudan Laboratory TODAY. DNR staff and an outside electrical contractor installed a new power cable between the 22nd and 27th levels. Lab staff then checked to be sure that all circuits in the lab were shut off. Following this check, power was turned on to the Lab substation via the new cable. Lab staff then turned on the lights and HVAC system in the Lab. These systems appear to function normally. The Lab closed loop cooling system also seems to function normally. The Lab fire detection system is also functioning and will shut down power to the lab, in the event of an alarm detection.

• The fiber optic cable has sufficient slack that it will be possible to cut out the damaged section and re-splice the cable. The splicing contractor will be working next week.

• Representatives of cleaning companies will visit the lab on Monday to prepare estimates for general clean up of the lab, including remnants of the foam and the debris brought in by the foam.

• ECI, the contractor hired by the DNR to do shaft renovations, will have key personnel on site this coming week, beginning with an on-site pre-construction meeting on Monday. CNA Engineering has also been hired to provide geo-technical advice to ECI. A CNA engineer will also be on site this coming week. A likely arrangement beginning this coming week is that lab and DNR access to the mine will occur on a first shift, possibly beginning at 6 am, with ECI shaft work on a second shift, possibly beginning at 2 pm.

• Work in the lab this coming week will focus on re-starting the MINOS Detector, beginning with checks of control system electronics and work to dry out the magnet cables.

• The sprinkler contractor has submitted drawings to the University Fire Marshal for the repairs to the sprinkler system. Approval of the drawings and repair work is also expected to start this coming week.

• The University has hired CNA Engineering and Don Zeni, a specialized drilling consultant, to evaluate the feasibility of providing an alternate access to the lab through either a ~6 inch diameter hole and/or a ~4 foot diameter hole. A small hole would enable an independent path for communications, power and high pressure air ventilation between the surface and the lab. The larger hole would also enable a secondary, emergency egress, in addition to the telecom and power connections.

planning.txt · Last modified: 2011/05/09 12:33 by 24.245.16.19